# Convergence Rates for the Maximum A Posteriori Estimator in PDE-Regression Models with Random Design

### v1

@@ Appendix A Appendix
In this section, we provide detailed proofs of the main results presented in Section 2. We start with the proof of Proposition 2.5 - the existence result for the MAP-estimator. To do so, fix $N\in\mathbb{N}$ as well as $\mathfrak{r}\in\mathbb{R}_{>0}$ throughout. The proof follows the approach of Nickl et al. [2020], where the existence of $\hat{\theta}_{\textsc{MAP}}$ is shown for the white noise formulation of the model. The authors make use of the direct method of the calculus of variations (see Dacorogna [2008] for instance) and split the proof into two parts: in one part it is shown that $\hat{\theta}_{\textsc{MAP}}$ exists as a maximizer of the functional $\mathscr{J}_{\mathfrak{r},N}$ over bounded subsets of $\Theta$ (local existence). In the other part it is shown that this already suffices to conclude the existence of $\hat{\theta}_{\textsc{MAP}}$ as a maximizer over the whole parameter space $\Theta$ (global existence). In the present work, we adapt the proof idea to the random design regression setting. We need the functional $\mathscr{J}_{\mathfrak{r},N}$ to be lower semicontinuous with respect to the weak topology of Sobolev spaces. Let $\tilde{\mathscr{H}}$ be either $H^{\alpha}(\mathcal{O})$ or $H^{\alpha}_{c}(\mathcal{O})$. Let $\mathcal{T}_{w}$ denote the weak topology of $\tilde{\mathscr{H}}$, i.e., the coarsest topology with respect to which all bounded linear functionals $L:\tilde{\mathscr{H}}\to\mathbb{R}$ are continuous. With a slight abuse of notation, the subspace topology on subsets of $\tilde{\mathscr{H}}$ is also denoted by $\mathcal{T}_{w}$. By standard theory of functional analysis (see, for instance, Megginson [1998], Theorem 2.6.23), on any closed ball ~$\tilde{\mathscr{H}}(r):=\{\theta\in\tilde{\mathscr{H}}:\,\left\|\theta\right\|_{H^{\alpha}(\mathcal{O})}\leq r\}$ of radius $r\in\mathbb{R}_{>0}$, the weak topology $\mathcal{T}_{w}$ is metrizable by some metric $\tilde{d}$. Defining analogously :=×$\mathscr{H}:=\bigtimes_{i\in\left\llbracket d_{W}\right\rrbracket}\tilde{\mathscr{H}}$, the weak topology $\mathcal{T}_{w}^{W}$ on $\mathscr{H}$ equals the product ×$\bigtimes_{i\in\left\llbracket d_{W}\right\rrbracket}\mathcal{T}_{w}$ (see proof of Theorem 3.10 in Brezis [2011]). By a standard fact of functional analysis, on each closed ball $\mathscr{H}(r):=\{\theta\in\mathscr{H}:\,\left\|\theta\right\|_{H^{\alpha}(\mathcal{O},W)}\leq r\}$, the weak topology $\mathcal{T}_{w}^{W}$ is then also metrizable by some metric $d$. In the following we claim that $-\mathscr{J}_{\mathfrak{r},N}:\Theta\cap\mathscr{H}(2^{j})\to\mathbb{R}\cup\{\infty\}$ is lower semicontinous with respect to $\mathcal{T}_{w}^{W}$ for any $j\in\mathbb{N}$. To prove this, we need the following general result from functional analysis.
Lemma § A.1 (Compact operators and weak convergence).
@@
Let $(X,\left\|\cdot\right\|_{X})$ and $(Y,\left\|\cdot\right\|_{Y})$ be two Banach spaces. Let $T:X\to Y$ be a linear compact operator. Assume that $(x_{n})_{n\in\mathbb{N}}\subseteq X$ is a weakly convergent sequence with limit $x\in X$. Then, $(T[x_{n}])_{n\in\mathbb{N}}\subseteq Y$ converges strongly in $Y$, i.e. $\lim_{n\to\infty}\left\|T[x_{n}]-T[x]\right\|_{Y}=0$.
Proof of 1 (Lemma A.1).
@@
Let $(x_{n})_{n\in\mathbb{N}}\subseteq X$ be a weakly convergent sequence with limit $x\in X$, i.e., for all $L\in X^{*}$, we have $\lim_{n\to\infty}L[x_{n}]=L[x]$. By an application of the uniform boundness principle, there exists $B\in\mathbb{R}_{>0}$, such that $\left\|x_{n}\right\|_{X}\leq B$ for all $n\in\mathbb{N}$. By hypothesis on $T$ being compact, the set $T((x_{n})_{n\in\mathbb{N}})$ is relative compact in $Y$, i.e. $\overline{T((x_{n})_{n\in\mathbb{N}})}$ is compact in $Y$. Hence, there exists a subsequence $(T[x_{n_{m}}])_{m\in\mathbb{N}}\subseteq Y$ with $\lim_{m\to\infty}\left\|T[x_{n_{m}}]-v\right\|_{Y}=0$ for some $v\in Y$. Further, for any $A\in Y^{*}$, we have that $A\circ T\in X^{*}$. Thus, $\lim_{n\to\infty}A[T[x_{n}]]=A[T[x]]$, such that the sequence $(T[x_{n}])_{n\in\mathbb{N}}\subseteq Y$ converges weakly in $Y$ with limit $T[x]\in Y$. By uniquess of the weak limit, we deduce $v=T[x]$. Overall, we have that the sequence $(T[x_{n_{m}}])_{m\in\mathbb{N}}\subseteq Y$ converges strongly in $Y$. Now, we see that each subsequence $(T[x_{n_{m_{l}}}])_{l\in\mathbb{N}}\subseteq Y$ converges strongly in $Y$, which already proves the claim.
Remark § A.2 (Embeddings and weak convergence).
@@
For $d\in\mathbb{N}$ let $\mathcal{O}\subseteq\mathbb{R}^{d}$ be a bounded domain with smooth boundary. Denoting by $B_{pq}^{s}(\mathcal{O})$ the ususal Besov space with parameters $p,q\in\mathbb{R}_{>0}\cup\{\infty\}$ and $s\in\mathbb{R}$, we have that the embedding
@@
is compact if $p_{1},p_{2},q_{1},q_{2}\in\mathbb{R}_{>0}\cup\{\infty\}$, $-\infty<s_{2}<s_{1}<\infty$ and $v:=s_{1}-s_{2}-d\max\{\frac{1}{p_{1}}-\frac{1}{p_{2}},0\}\in\mathbb{R}_{>0}$ (see Edmunds and Triebel [1996] for definitions and properties). Let $\alpha\in\mathbb{R}_{>0}$ and $\eta\in\mathbb{R}_{\geq 0}\setminus\mathbb{N}$. Identifying $H^{\alpha}(\mathcal{O})=B_{22}^{\alpha}(\mathcal{O})$, $\mathbb{L}^{2}(\mathcal{O})=B_{22}^{0}(\mathcal{O})$ and $C^{\eta}(\mathcal{O})=B_{\infty\infty}^{\eta}(\mathcal{O})$, we have in particular, that the embeddings $H^{\alpha}(\mathcal{O})\hookrightarrow\mathbb{L}^{2}(\mathcal{O})$ and $H^{\alpha}(\mathcal{O})\hookrightarrow C^{\eta}(\mathcal{O})$ for $\alpha>\eta+\frac{d}{2}$ are compact. Further, following Proposition 4.5 in Triebel [2006], we have for any $k\in\mathbb{N}$ the continous embedding $B_{\infty 1}^{k}(\mathcal{O})\hookrightarrow C^{k}(\overline{\mathcal{O}})$. Making use of above results, we thus have that the embedding $H^{\alpha}(\mathcal{O})\hookrightarrow C^{k}(\mathcal{O})$ is also compact, provided that $\alpha>k+\frac{d}{2}$. If $(W,\left|\cdot\right|_{W})$ denotes a $d_{W}$-dimensional $\mathbb{R}$-vector space, the same results apply, if we replace the function spaces $F(\mathcal{O})$ by their tensor products $F(\mathcal{O},W)$. Now let $(\theta_{n})_{n\in\mathbb{N}}\subseteq\mathscr{H}$ be a sequence, such that $\theta_{n}\to\theta\in\mathscr{H}$ with respect to $\mathcal{T}_{w}^{W}$. Then, an application of Lemma A.1 yields that $(\theta_{n})_{n\in\mathbb{N}}$ converges strongly in $\mathbb{L}^{2}(\mathcal{O},W)$ as well as in $C^{\eta}(\mathcal{O},W)$, if $\alpha>\eta+\frac{d}{2}$ for $\eta\in\mathbb{R}_{\geq 0}$.
We derive that the forward map $\mathscr{G}$ is continuous, which is necessary to prove the lower semicontinuity of $-\mathscr{J}_{\mathfrak{r},N}$.
Lemma § A.3 (Continuity of $\mathscr{G}$).
For $\gamma_{2},\eta_{1},\eta_{2}\in\mathbb{R}_{\geq 0}$ let $\mathscr{G}:\Theta\to\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)$ satisfy C3) of 2.3. Let $\Theta_{o}\subseteq\Theta$ be a bounded subset of $\mathscr{H}$ with $\alpha>\eta_{1}\lor\eta_{2}+\frac{d}{2}$. Let $(\Omega,\mathcal{A},\mathbb{P})$ be the underlying probability space of $Z:\Omega\to\mathcal{Z}$. Then, the map
$(\Theta_{o},d)\ni\theta\mapsto\mathscr{G}(\theta)(Z)\in(V,\left|\cdot\right|_{V})$
@@
Proof of 2 (Lemma A.3).
Fix some $\omega\in\Omega$ (outside of a $\mathbb{P}$-null-set). Let $(\theta_{n})_{n\in\mathbb{N}}\subseteq\Theta_{o}$ be a convergent sequence with weak limit $\theta\in\Theta_{o}$, i.e., $\lim_{n\to\infty}d(\theta_{n},\theta)=0$. Note that $\left\|\theta_{n}\right\|_{H^{\alpha}(\mathcal{O},W)}\lor\left\|\theta\right\|_{H^{\alpha}(\mathcal{O},W)}\leq B$ for some $B\in\mathbb{R}_{>0}$ uniformly over all $n\in\mathbb{N}$. The claim is proven, if $(\varphi_{n}(Z(\omega)))_{n\in\mathbb{N}}\subseteq\mathbb{R}$, defined by
$\varphi_{n}(Z(\omega)):=\left|\mathscr{G}(\theta_{n})(Z(\omega))-\mathscr{G}(\theta)(Z(\omega))\right|_{V}$
is a null-sequence. By hypothesis C3) and the continuous embedding $H^{\alpha}(\mathcal{O},W)\hookrightarrow C^{\eta_{1}}(\mathcal{O},W)$, there exists a universal constant $c\in\mathbb{R}_{>0}$, such that
$\varphi_{n}(Z(\omega))\leq\left\|\mathscr{G}(\theta_{n})-\mathscr{G}(\theta)\right\|_{\infty}$ $\leq c(1+B^{\gamma_{2}})\left\|\theta_{n}-\theta\right\|_{C^{\eta_{2}}(\mathcal{O},W)}.$
Exploiting Remark A.2, the right hand side converges to zero, which proves the claim.
Lemma § A.4 (Uniform continuity of $\mathscr{G}$).
Let $\mathscr{G}$ satisfy C2) and C3) of 2.3. Let $D_{N}=((Y_{i},Z_{i}))_{i\in\left\llbracket N\right\rrbracket}\sim\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}$ for $N\in\mathbb{N}$. Let $r\in\mathbb{R}_{>0}$. Then, the map
$E:(\Theta\cap\mathscr{H}(r),d)\to(\mathbb{R},|\cdot|),\,\theta\mapsto E[\theta]:=\frac{1}{2N\sigma^{2}}\sum_{i\in\left\llbracket N\right\rrbracket}\left|Y_{i}-\mathscr{G}(\theta)(Z_{i})\right|_{V}^{2}$
@@
Proof of 3 (Lemma A.4).
For any $\delta\in\mathbb{R}_{>0}$, let $M_{\delta}$ be the modulus of continuity of $E$, i.e.,
$M_{\delta}:=\sup_{\theta,\tilde{\theta}\in\Theta\cap\mathscr{H}(r):d(\theta,\tilde{\theta})\leq\delta}\left\{|E[\theta]-E[\tilde{\theta}]|\right\},$
@@
$A:=\left\{\omega\in\Omega:\,\lim_{\delta\to 0}M_{\delta}=0\right\}.$
@@
$|E[\theta]-E[\tilde{\theta}]|$ $\leq\frac{2C_{\mathscr{G},\operatorname{B}}}{N\sigma^{2}}\sum_{i\in\left\llbracket N\right\rrbracket}\left|\mathscr{G}(\tilde{\theta})(Z_{i})-\mathscr{G}(\theta)(Z_{i})\right|_{V}+\frac{1}{N\sigma}\sum_{i\in\left\llbracket N\right\rrbracket}|\langle\varepsilon_{i},\mathscr{G}(\tilde{\theta})(Z_{i})-\mathscr{G}(\theta)(Z_{i})\rangle_{V}|.$
Analogously, define for each $\delta\in\mathbb{R}_{>0}$ and $Z\sim\zeta$ the modulus of continuity of $\mathscr{G}$ by
$R_{\delta}:=\sup_{\theta,\tilde{\theta}\in\Theta\cap\mathscr{H}(r):d(\theta,\tilde{\theta})\leq\delta}\left\{\left|\mathscr{G}(\tilde{\theta})(Z)-\mathscr{G}(\theta)(Z)\right|_{V}\right\},$
which in turns defines a random variable. In Lemma A.3, we have shown that $\mathscr{G}$ is almost sure continuous as a mapping from $\Theta\cap\mathscr{H}(r)$ to $V$. As $\Theta\cap\mathscr{H}(r)$ is a compact metric space, $\mathscr{G}$ is in fact almost surely uniformly continuous, and hence $\lim_{\delta\to 0}R_{\delta}=0$ almost surely. Moreover, since $Z$ and $\varepsilon\sim N(0,\operatorname{Id}_{V})$ are independent by hypothesis, we have
$\mathbb{E}\left[\sup_{\theta,\tilde{\theta}\in\Theta\cap\mathscr{H}(r):d(\theta,\tilde{\theta})\leq\delta}\left\{|\langle\varepsilon,\mathscr{G}(\tilde{\theta})(Z)-\mathscr{G}(\theta)(Z)\rangle_{V}|\right\}\right]$ $\leq\mathbb{E}[\left|\varepsilon\right|_{V}R_{\delta}]=\sqrt{d_{V}}\cdot\mathbb{E}[R_{\delta}].$
@@
which proves the claim.
We are now providing the proof of Proposition 2.5.
Proof of 4 (Proposition 2.5).
Grant the assumptions of Proposition 2.5. We follow along the lines of Nickl et al. [2020] and split the proof of existence in two parts.
@@
    Local Existence: Let $j\in\mathbb{N}$ be arbitrarily. We show now that $\mathscr{J}_{\mathfrak{r},N}$ has a maximizer over $\Theta\cap\mathscr{H}(2^{j})$. As $\Theta$ is assumed to be weakly closed and $\mathscr{H}(2^{j})$ is weakly sequentially compact, due to the separability and reflexivity of Sobolev spaces and by an applycation of the Banach-Alaoglu theorem, any sequence $(\theta_{n})_{n\in\mathbb{N}}\subseteq\Theta\cap\mathscr{H}(2^{j})$ has a weakly convergent sub-sequence $(\theta_{n_{k}})_{k\in\mathbb{N}}$ with weak limit $\theta\in\Theta\cap\mathscr{H}(2^{j})$. Thus, following the Direct Method of the Calculus of Variations, it suffices to show that $-\mathscr{J}_{\mathfrak{r},N}|_{\Theta\cap\mathscr{H}(2^{j})}$ is almost surely weakly lower semicontinuous. Rewrite $-\mathscr{J}_{\mathfrak{r},N}|_{\Theta\cap\mathscr{H}(2^{j})}$ as
    $-\mathscr{J}_{\mathfrak{r},N}|_{\Theta\cap\mathscr{H}(2^{j})}(\theta)=E_{1}(\theta)+E_{2}(\theta)$ (A.1)
@@
    $E_{1}(\theta):=\frac{1}{2N\sigma^{2}}\sum_{i\in\left\llbracket N\right\rrbracket}\left|Y_{i}-\mathscr{G}(\theta)(Z_{i})\right|_{V}^{2}\text{ and }E_{2}(\theta):=\frac{\mathfrak{r}^{2}}{2}\left\|\theta\right\|_{H^{\alpha}(\mathcal{O},W)}^{2}.$
    $E_{2}$ is weakly lower semicontinuous, which is a standard result in functional analysis concerning norms on Banach spaces. In Lemma A.4, we have derived that $E_{1}$ is almost surely uniformly continuous as a map between the metric spaces $(\Theta\cap\mathscr{H}(2^{j}),d)$ and $(\mathbb{R},|\cdot|)$. In particular, $E_{1}$ is weakly lower semicontinuous. All in all, local existence of $\hat{\theta}_{\textsc{MAP}}$ follows immediately by an application of the Direct Methods of Calculus of Variations (see Dacorogna [2008]).
@@
    Then, $(A_{j})_{j\in\mathbb{N}}$ satisfies $\lim_{j\to\infty}\mathbb{P}[A_{j}]=0$ according to Theorem 2.6. Thus, choosing $j\in\mathbb{N}$ large enough, we have
    $\sup_{\theta\in\Theta\cap\mathscr{H}(2^{j})}\left\{\mathscr{J}_{\mathfrak{r},N}[F]\right\}=\sup_{\theta\in\Theta}\left\{\mathscr{J}_{\mathfrak{r},N}[F]\right\}$

### v3

@@ 4.1 Proofs of section 2
In this section, we provide detailed proofs of the main results presented in section 2. We start with the proof of proposition 2.5 - the existence result for the MAP-estimator. To do so, fix $N\in\mathbb{N}$ as well as $\mathfrak{r}\in\mathbb{R}_{>0}$ throughout. The proof follows the approach of Nickl et al. (2020), where the existence of $\hat{\theta}_{\textsc{MAP}}$ is shown for the White Noise formulation of the model. The authors make use of the direct method of the calculus of variations (see Dacorogna (2008) for instance) and split the proof into two parts: Firstly, it is shown that $\hat{\theta}_{\textsc{MAP}}$ exists as a maximizer of the functional $\mathscr{J}_{\mathfrak{r},N}$ over bounded subsets of $\Theta$ (local existence). Secondly, it is shown that this already suffices to conclude the existence of $\hat{\theta}_{\textsc{MAP}}$ as a maximizer over the whole parameter space $\Theta$ (global existence). In the present work, we adapt the proof idea to the random design regression setting. We need the functional $\mathscr{J}_{\mathfrak{r},N}$ to be lower semicontinuous with respect to the weak topology of Sobolev spaces.
Let $\tilde{\mathscr{H}}$ be either $H^{\alpha}(\mathcal{O})$ or $H^{\alpha}_{c}(\mathcal{O})$. Let $\mathcal{T}_{w}$ denote the weak topology of $\tilde{\mathscr{H}}$, i.e., the coarsest topology with respect to which all bounded linear functionals $L:\tilde{\mathscr{H}}\to\mathbb{R}$ are continuous. With a slight abuse of notation, the corresponding subspace topology on subsets of $\tilde{\mathscr{H}}$ is also denoted by $\mathcal{T}_{w}$. By standard theory of functional analysis (see for instance Megginson (1998), Theorem 2.6.23), on any closed ball $\tilde{\mathscr{H}}(r):=\{\theta\in\tilde{\mathscr{H}}:\,\|\theta\|_{H^{\alpha}(\mathcal{O})}\leq r\}$ of radius $r\in\mathbb{R}_{>0}$, the weak topology $\mathcal{T}_{w}$ is metrizable by some metric $\tilde{d}$. Defining the product space $\mathscr{H}:=\bigtimes_{i\in\left\llbracket d_{W}\right\rrbracket}\tilde{\mathscr{H}}$, the weak topology $\mathcal{T}_{w}^{W}$ on $\mathscr{H}$ equals to $\bigtimes_{i\in\left\llbracket d_{W}\right\rrbracket}\mathcal{T}_{w}$ (see proof of Theorem 3.10 in Brezis (2011)). As above, on each closed ball $\mathscr{H}(r):=\{\theta\in\mathscr{H}:\,\|\theta\|_{H^{\alpha}(\mathcal{O},W)}\leq r\}$, the weak topology $\mathcal{T}_{w}^{W}$ is then also metrizable by some metric $d$.
x
In the following we claim that $\mathscr{J}_{\mathfrak{r},N}:\Theta\cap\mathscr{H}(2^{j})\to\mathbb{R}\cup\{\infty\}$ is lower semicontinous with respect to $\mathcal{T}_{w}^{W}$ for any $j\in\mathbb{N}$. To this end, we need the following Lemma.
Lemma § 4.1 (Compact Operators and Weak Convergence).
@@
Let $(X,\|\cdot\|_{X})$ and $(Y,\|\cdot\|_{Y})$ be two Banach spaces. Let $T:X\to Y$ be a linear and compact operator. Assume that $(x_{n})_{n\in\mathbb{N}}\subseteq X$ is a weakly convergent sequence with limit $x\in X$. Then, $(T[x_{n}])_{n\in\mathbb{N}}\subseteq Y$ converges strongly in $Y$, i.e. $\lim_{n\to\infty}\|T[x_{n}]-T[x]\|_{Y}=0$.
Proof of lemma 4.1.
@@
Let $(x_{n})_{n\in\mathbb{N}}\subseteq X$ be a weakly convergent sequence with limit $x\in X$, i.e., for all $L\in X^{*}$, it holds $\lim_{n\to\infty}L[x_{n}]=L[x]$. By an application of the uniform boundedness principle, there exists $B\in\mathbb{R}_{>0}$, such that $\|x_{n}\|_{X}\leq B$ for all $n\in\mathbb{N}$. In other words, $(x_{n})_{n\in\mathbb{N}}$ is bounded as well as every sub-sequence of $(x_{n})_{n\in\mathbb{N}}$. By hypothesis on $T$ being compact, the set $T((x_{n})_{n\in\mathbb{N}})$ is relative compact in $Y$, i.e. $\overline{T((x_{n})_{n\in\mathbb{N}})}$ is compact in $Y$. Thus, any subsequence $(T[x_{n_{m}}])_{m\in\mathbb{N}}$ of $(T[x_{n}])_{n\in\mathbb{N}}$ has a further strongly converging subsequence $(T[x_{n_{m_{l}}}])_{l\in\mathbb{N}}$ with limit $y\in Y$, that is $\lim_{l\to\infty}\|T[x_{n_{m_{l}}}]-y\|_{Y}=0$. Further, observe that $(T[x_{n}])_{n\in\mathbb{N}}$ converges weakly in $Y$, as for any $A\in Y^{*}$ we have $A\circ T\in X^{*}$ and thus $\lim_{n\to\infty}A[T[x_{n}]]=\lim_{n\to\infty}(A\circ T)[x_{n}]=(A\circ T)[x]=A[T[x]]$. In other words, we have shown that $(T[x_{n}])_{n\in\mathbb{N}}$ converges weakly to $T[x]$ in $Y$ and $(T[x_{n_{m_{l}}}])_{l\in\mathbb{N}}$ converges strongly to $y$ in $Y$. By uniqueness of the weak limit, we deduce $y=T[x]$ and therefore $T[x_{n}]$ converges strongly to $T[x]$ in $Y$.
Remark § 4.2 (Embeddings and Weak Convergence).
@@
For $d\in\mathbb{N}$ let $\mathcal{O}\subseteq\mathbb{R}^{d}$ be a bounded domain with smooth boundary. Denoting by $B_{pq}^{s}(\mathcal{O})$ the ususal Besov space with parameters $p,q\in\mathbb{R}_{>0}\cup\{\infty\}$ and $s\in\mathbb{R}$, the embedding
@@
is compact, if $p_{1},p_{2},q_{1},q_{2}\in\mathbb{R}_{>0}\cup\{\infty\}$, $\infty<s_{2}<s_{1}<\infty$ and $s_{1}-s_{2}-d\max\{\frac{1}{p_{1}}-\frac{1}{p_{2}},0\}\in\mathbb{R}_{>0}$ (see Edmunds and Triebel (1996) for definitions and properties). Let $\alpha\in\mathbb{R}_{>0}$ and $\eta\in\mathbb{R}_{\geq 0}\setminus\mathbb{N}$. In particular, by identifying $H^{\alpha}(\mathcal{O})=B_{22}^{\alpha}(\mathcal{O})$, $\mathbb{L}^{2}(\mathcal{O})=B_{22}^{0}(\mathcal{O})$ and $C^{\eta}(\mathcal{O})=B_{\infty\infty}^{\eta}(\mathcal{O})$, the embeddings $H^{\alpha}(\mathcal{O})\hookrightarrow\mathbb{L}^{2}(\mathcal{O})$ and $H^{\alpha}(\mathcal{O})\hookrightarrow C^{\eta}(\mathcal{O})$ for $\alpha>\eta+\frac{d}{2}$ are compact. Further, following Proposition 4.5 in Triebel (2006), for any $k\in\mathbb{N}$ the embedding $B_{\infty 1}^{k}(\mathcal{O})\hookrightarrow C^{k}(\overline{\mathcal{O}})$ is continuous. Hence, the embedding $H^{\alpha}(\mathcal{O})\hookrightarrow C^{k}(\mathcal{O})$ is also compact, provided that $\alpha>k+\frac{d}{2}$. If $(W,\left|\cdot\right|_{W})$ denotes a $d_{W}$-dimensional $\mathbb{R}$-vector space, the same results apply, if we replace the function spaces $F(\mathcal{O})$ by their tensor products $F(\mathcal{O},W)$. With $\mathscr{H}$ as defined above, let $(\theta_{n})_{n\in\mathbb{N}}\subseteq\mathscr{H}$ be a sequence, such that $\theta_{n}\to\theta\in\mathscr{H}$ with respect to $\mathcal{T}_{w}^{W}$. An application of lemma 4.1 then yields that $(\theta_{n})_{n\in\mathbb{N}}$ converges strongly in $\mathbb{L}^{2}(\mathcal{O},W)$ as well as in $C^{\eta}(\mathcal{O},W)$, if $\alpha>\eta+\frac{d}{2}$ for $\eta\in\mathbb{R}_{\geq 0}$.
Lemma § 4.3 (Continuity of $\mathscr{G}$).
For $\gamma_{2},\eta_{1},\eta_{2}\in\mathbb{R}_{\geq 0}$, let $\mathscr{G}:\Theta\to\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)$ satisfy C3) of 2.3. Let $\Theta_{o}\subseteq\Theta$ be a bounded subset of $\mathscr{H}$ with $\alpha>\eta_{1}\lor\eta_{2}+\frac{d}{2}$. Let $(\Omega,\mathcal{A},\mathbb{P})$ be the underlying probability space of $Z:\Omega\to\mathcal{Z}$. Then, the map
$(\Theta_{o},d)\ni\theta\mapsto\mathscr{G}(\theta)(Z)\in(V,|\cdot|_{V})$
@@
Proof of lemma 4.3.
Fix some $\omega\in\Omega$ (outside of a $\mathbb{P}$-null-set). Let $(\theta_{n})_{n\in\mathbb{N}}\subseteq\Theta_{o}$ be a convergent sequence with weak limit $\theta\in\Theta_{o}$, i.e., $\lim_{n\to\infty}d(\theta_{n},\theta)=0$. Note that $\|\theta_{n}\|_{H^{\alpha}(\mathcal{O},W)}\lor\|\theta\|_{H^{\alpha}(\mathcal{O},W)}\leq B$ for some $B\in\mathbb{R}_{>0}$ uniformly over all $n\in\mathbb{N}$. The claim is proven, if $(\varphi_{n}(Z(\omega)))_{n\in\mathbb{N}}\subseteq\mathbb{R}$, defined by
$\varphi_{n}(Z(\omega)):=|\mathscr{G}(\theta_{n})(Z(\omega))-\mathscr{G}(\theta)(Z(\omega))|_{V}$
is a null-sequence. By hypothesis and exploiting the continuous embedding $H^{\alpha}(\mathcal{O},W)\hookrightarrow C^{\eta_{1}}(\mathcal{O},W)$, there exists a constant $c\in\mathbb{R}_{>0}$ (depending on $\alpha,\eta_{1},\eta_{2},d,C_{\operatorname{Lip},\infty}$), such that
$\varphi_{n}(Z(\omega))\leq\|\mathscr{G}(\theta_{n})-\mathscr{G}(\theta)\|_{\infty}$ $\leq c\cdot(1+B^{\gamma_{2}})\cdot\|\theta_{n}-\theta\|_{C^{\eta_{2}}(\mathcal{O},W)}.$
Exploiting remark 4.2, we have $\lim_{n\to\infty}\|\theta_{n}-\theta\|_{C^{\eta_{2}}(\mathcal{O},W)=0}$, which proves the claim. ∎
Lemma § 4.4 (Uniform Continuity).
Let $\mathscr{G}:\Theta\to\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)$ satisfy C2) and C3) of 2.3. Let $D_{N}=((Y_{i},Z_{i}))_{i\in\left\llbracket N\right\rrbracket}\sim\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}$ for $N\in\mathbb{N}$ as in theorem 2.6. Let $r\in\mathbb{R}_{>0}$. Then, the map
$E:(\Theta\cap\mathscr{H}(r),d)\to(\mathbb{R},|\cdot|),\,\theta\mapsto E[\theta]:=\frac{1}{2N\sigma^{2}}\sum_{i\in\left\llbracket N\right\rrbracket}|Y_{i}-\mathscr{G}(\theta)(Z_{i})|_{V}^{2}$
@@
Proof of lemma 4.4.
For any $\delta\in\mathbb{R}_{>0}$, denote by $M_{\delta}$ the modulus of continuity of $E$, i.e.,
[…]
@@
$A:=\left\{\omega\in\Omega:\,\lim_{\delta\to 0}M_{\delta}(\omega)=0\right\}.$
@@
[…]
Define for each $\delta\in\mathbb{R}_{>0}$ and $Z\sim\zeta$ the modulus of continuity of $\mathscr{G}$ by
$R_{\delta}:=\sup_{\theta,\tilde{\theta}\in\Theta\cap\mathscr{H}(r):d(\theta,\tilde{\theta})\leq\delta}\left\{|\mathscr{G}(\tilde{\theta})(Z)-\mathscr{G}(\theta)(Z)|_{V}\right\},$
which in turn defines also a random variable. In lemma 4.3, we have shown that $\mathscr{G}$ is almost surely continuous as a mapping from $\Theta\cap\mathscr{H}(r)$ to $V$. As $\Theta\cap\mathscr{H}(r)$ is a compact metric space, $\mathscr{G}$ is in fact almost surely uniformly continuous, and hence $\lim_{\delta\to 0}R_{\delta}=0$ almost surely. Moreover, since $Z$ and $\varepsilon\sim N(0,\operatorname{Id}_{V})$ are independent by hypothesis, we have
$\mathbb{E}\left[\sup_{\theta,\tilde{\theta}\in\Theta\cap\mathscr{H}(r):d(\theta,\tilde{\theta})\leq\delta}\left\{|\langle\varepsilon,\mathscr{G}(\tilde{\theta})(Z)-\mathscr{G}(\theta)(Z)\rangle_{V}|\right\}\right]$ $\leq\mathbb{E}[|\varepsilon|_{V}R_{\delta}]=\sqrt{d_{V}}\cdot\mathbb{E}[R_{\delta}].$
@@
which shows the claim. ∎
Proof of proposition 2.5.
Grant the assumptions of proposition 2.5. We follow along the lines of Nickl et al. (2020) and split the proof of existence in two parts.
@@
    Local Existence: Let $j\in\mathbb{N}$ be arbitrarily. We show now that $\mathscr{J}_{\mathfrak{r},N}$ has a maximizer over $\Theta\cap\mathscr{H}(2^{j})$. As $\Theta$ is assumed to be weakly closed and $\mathscr{H}(2^{j})$ is weakly sequentially compact, due to the separability and reflexivity of Sobolev spaces and by an application of the Banach-Alaoglu theorem, any sequence $(\theta_{n})_{n\in\mathbb{N}}\subseteq\Theta\cap\mathscr{H}(2^{j})$ has a weakly convergent sub-sequence $(\theta_{n_{k}})_{k\in\mathbb{N}}$ with weak limit $\theta\in\Theta\cap\mathscr{H}(2^{j})$. Thus, following the Direct Method of the Calculus of Variations, it suffices to show that $\mathscr{J}_{\mathfrak{r},N}|_{\Theta\cap\mathscr{H}(2^{j})}$ is almost surely weakly lower semicontinuous. Rewrite $\mathscr{J}_{\mathfrak{r},N}|_{\Theta\cap\mathscr{H}(2^{j})}$ as
    $\mathscr{J}_{\mathfrak{r},N}|_{\Theta\cap\mathscr{H}(2^{j})}(\theta)=E_{1}(\theta)+E_{2}(\theta)$ (17)
@@
    $E_{1}(\theta):=\frac{1}{2N\sigma^{2}}\sum_{i\in\left\llbracket N\right\rrbracket}|Y_{i}-\mathscr{G}(\theta)(Z_{i})|_{V}^{2}\text{ and }E_{2}(\theta):=\frac{\mathfrak{r}^{2}}{2}\|\theta\|_{H^{\alpha}(\mathcal{O},W)}^{2}.$
    $E_{2}$ is weakly lower semicontinuous, which is a standard result in functional analysis concerning norms on Banach spaces. In lemma 4.4, we have derived that $E_{1}$ is almost surely uniformly continuous as a map between the metric spaces $(\Theta\cap\mathscr{H}(2^{j}),d)$ and $(\mathbb{R},|\cdot|)$. In particular, $E_{1}$ is weakly lower semicontinuous. All in all, local existence of $\hat{\theta}_{\textsc{MAP}}$ follows immediately by an application of the Direct Methods of Calculus of Variations (see Dacorogna (2008)).
@@
    Then, $(A_{j})_{j\in\mathbb{N}}$ satisfies $\lim_{j\to\infty}\mathbb{P}[A_{j}]=0$ according to the concentration inequality in theorem 2.6 i). Thus, choosing $j\in\mathbb{N}$ large enough, we have
    $\sup_{\theta\in\Theta\cap\mathscr{H}(2^{j})}\left\{\mathscr{J}_{\mathfrak{r},N}[\theta]\right\}=\sup_{\theta\in\Theta}\left\{\mathscr{J}_{\mathfrak{r},N}[\theta]\right\}$